### v1

@@ 3 Large-Scale FEM Data
For Kelvinlet regularization, interaction points are sampled uniformly from the surface nodes. Initial displacements are then drawn uniformly from a cone centered around the surface normal at $\mathbf{x}_{s}$, with a maximum opening angle of $\alpha_{\max}$.
4 Texture Rendering for Realistic Simulations
Our visualization pipeline combines real-time tissue deformation and high-fidelity rendering to create a realistic simulation (see Figure 1 and the supplementary videos). In a preprocessing stage, surface meshes are extracted from the segmented high-resolution volumetric mesh, assigned texture coordinates, and enriched with procedurally generated textures using Worley noise [26]. Interpolation weights are computed to map low-resolution deformations to the high-resolution mesh [16], optimized using a Bounding Volume Hierarchy (BVH) [5], and UV coordinates are transferred to the volumetric mesh. At runtime, the system dynamically updates surface meshes to accommodate added or removed tetrahedral elements, interpolates high-resolution vertex positions based on low-resolution deformations, smooths exposed internal structures, and recalculates surface normals. Rendering employs pre-integrated diffuse shading [19], triplanar texture projection for inner tissue surfaces [7], and post-processing effects within Unity to achieve realistic tissue appearance. The pipeline achieves near-real-time performance while preserving visual fidelity during dynamic deformations.
@@ 5 Experiments
We evaluate Kelvinlet-based priors using four neural architectures: U-Net [15], SAB [13] (set attention), GraphGPS [23] (a Graph Neural Network and Transformer hybrid), and PointNet [21] (a universal function approximator for sets). We modify PointNet and GraphGPS by reducing layers, U-Net by replacing convolutions with shared MLPs to form a DeepSet, and SAB by using two layers with an MLP regressor head.
Each model is trained separately on two datasets: one with individual graspers and another with multiple graspers interacting. For each dataset, models are trained under three settings: (1) a baseline without physics priors (i.e., evaluating solely the network), (2) residual learning predicting deviations from Kelvinlets, and (3) physics-informed regularization enforcing Kelvinlet constraints.
@@ 5.1 Results and Analysis
Figure 2: Qualitative comparison of deformation predictions for the best (PointNet) and worst (GraphGPS) models in linear (top) and nonlinear (bottom) regimes.
To quantitatively assess model performance, we introduce the Deformation Capture Mean (DCM) metric, which measures how well a predicted deformation field approximates the ground-truth displacement field. The DCM is computed as:
@@
Tables 1 and 2 report DCMs across models and settings. Each experiment runs with three random seeds, reporting mean and standard deviation. Models perform better on linear cases, as nonlinear deformations introduce gravity effects, non-proportional stress-strain relationships, and greater sensitivity to perturbations. Notably, Kelvinlet priors consistently improve DCMs, despite being derived for linear elasticity, showing their effectiveness even in nonlinear regimes.
Table 1: DCM of models with/without Kelvinlet priors in linear simulations.
Model Indiv. Graspers Comb. Graspers
Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
U-Net[15] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08)
GraphGPS[23] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04)
PointNet[21] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07)
SAB[13] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (011) 91.02 (0.04)
Kelvinlet 65.24 (0) 64.33 (0)
Table 2: DCM of models with/without Kelvinlet priors in nonlinear simulations.
Model Indiv. Graspers Comb. Graspers
Base +Kelv. Res. +Kelv. Reg. Base +Kelv. Res. +Kelv. Reg.
U-Net[15] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21)
GraphGPS[23] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16)
PointNet[21] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11)
SAB[13] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16)
Kelvinlet 58.88 (0) 51.67 (0)
@@ 6 Conclusion
Acknowledgment
This work was partially supported by the NSF CAREER Award #2339898 and the Wellcome Leap ‘Surgery: Assess/Validate/Expand (SAVE)’ program.

### v2

@@ 2 Large-Scale FEM Data
For Kelvinlet regularization, interaction points $\mathbf{x}_{s}$ are sampled uniformly from the surface nodes. The corresponding displacements $\mathbf{u}_{s}\in\mathbb{R}^{3}$ are then sampled uniformly from a cone centered around the surface normal at $\mathbf{x}_{s}$, with an opening angle of $\alpha_{\max}$.
Having detailed the data generation process, we now turn to our proposed neural augmented Kelvinlet solution.
3 Neural-Augmented Kelvinlet
In this section, we present the Kelvinlets prior and our residual learning and regularization frameworks.
3.1 Navier-Cauchy Eqs and Green’s Function Solution
Figure 2: Overview of the two proposed training frameworks. Method 1 (Left): Residual learning predicts only the deviation from the Kelvinlet solution, reducing learning complexity by leveraging the physics-based analytical solution as an initialization. Method 2 (Right): Kelvinlet-based regularization combines FEM-supervised learning with a secondary loss enforcing consistency with Kelvinlet priors. The heatmap depicts the displacement norms.
where $\mu>0$ is the shear modulus, and $\nu\in(0,0.5)$ is the Poisson’s ratio, ensuring a physically consistent material response. Here, $\nabla$ denotes the gradient operator, $\nabla\cdot\mathbf{u}$ is the divergence of the displacement field, and $\nabla^{2}\mathbf{u}=\nabla\cdot(\nabla\mathbf{u})$ is the vector Laplacian applied to $\mathbf{u}$.
For a concentrated force $\mathbf{f}\delta(\mathbf{x}-\mathbf{x}_{s})$ applied at a source point $\mathbf{x}_{s}\in\mathbb{R}^{3}$, the Green’s function solution, known as the Kelvinlet, provides an analytical displacement field $\mathbf{u}:\mathbb{R}^{3}\setminus\{\mathbf{x}_{s}\}\to\mathbb{R}^{3}$ following:
$\begin{split}\mathbf{u}(\mathbf{x})=\frac{1}{4\pi\mu}\Bigg(\frac{(1-\nu)I}{r}+\nu\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r^{3}}\Bigg)\mathbf{f}\end{split}$ (6)
where $r=\|\mathbf{x}-\mathbf{x}_{s}\|_{2}$ denotes the Euclidean distance from the source, and $\otimes$ denotes the outer product operator.
$\mathbf{u}_{\varepsilon}(\mathbf{x})=\frac{\varepsilon}{(5-6\nu)}\Bigg(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{s})\otimes(\mathbf{x}-\mathbf{x}_{s})}{r_{\varepsilon}^{3}}\Bigg)\mathbf{u}_{s}.$ (7)
For multiple interaction points, we note that at each interaction point we only observe the superposition of contributions from all interaction points, without knowing the exact contribution of each individual point to this superposition. Consequently, unlike the single-grasper setting, $\mathbf{u}_{s}$ in Eq. (7) is not directly available for each interaction point and must be estimated. To formalize this, let $\mathbf{x}_{i}$ denote the location of the $i$th interaction, $\mathbf{u}_{i}$ the corresponding total displacement at $\mathbf{x}_{i}$, and $\mathbf{k}_{i}$ the unknown contribution of the $i$th interaction to this displacement. Assuming there are $K$ interaction points, the total displacement can be expressed as:
$\mathbf{u}_{\epsilon}(\mathbf{x})=\sum_{i=1}^{K}\underbrace{\frac{\varepsilon}{(5-6\nu)}\left(\frac{(3-4\nu)I}{r_{\varepsilon}}+\frac{(\mathbf{x}-\mathbf{x}_{i})\otimes(\mathbf{x}-\mathbf{x}_{i})}{r_{\varepsilon}^{3}}\right)}_{\Gamma_{i}(\mathbf{x})}\mathbf{k_{i}},$ (8)
where $\Gamma_{i}(\mathbf{x})\in\mathbb{R}^{3\times 3}$. Note that this formulation enables the total displacement function to incorporate cross communication between interaction points.
To estimate the $\mathbf{k}_{i}$ values, we solve the following optimization problem, which enforces the boundary conditions $\mathbf{u}_{\varepsilon}(\mathbf{x}_{i})=\mathbf{u}_{i}$ for all $i$:
$\operatorname*{argmin}_{\mathbf{k}}\sum_{i=1}^{K}\left(\|\sum_{j=1}^{K}\Gamma_{j}(\mathbf{x}_{i})\mathbf{k}_{j}-\mathbf{u}_{i}\|^{2})+\lambda\|\mathbf{k}_{i}\|^{2}\right).$ (9)
This formulation ensures that the individual Kelvinlet contributions collectively satisfy the prescribed displacements, while the regularization term $\lambda=0.001$ promotes a minimum-energy solution. The optimization in Eq. (9) is essentially a regularized least squares problem, which admits a fast analytic solution:
$\mathbf{k}^{*}=\left(\mathbf{\Gamma}^{\top}\mathbf{\Gamma}+\lambda\mathbf{I}\right)^{-1}\mathbf{\Gamma}^{\top}\mathbf{u},$ (10)
where $\mathbf{I}\in\mathbb{R}^{3K\times 3K}$ is the identity matrix, and $\mathbf{\Gamma}\in\mathbb{R}^{3K\times 3K}$, $\mathbf{k}\in\mathbb{R}^{3K}$, and $\mathbf{u}\in\mathbb{R}^{3K}$ are defined as:
$\mathbf{\Gamma}=\begin{bmatrix}\Gamma_{1}(\mathbf{x}_{1})&\cdots&\Gamma_{K}(\mathbf{x}_{1})\\ \vdots&\ddots&\vdots\\ \Gamma_{1}(\mathbf{x}_{K})&\cdots&\Gamma_{K}(\mathbf{x}_{K})\end{bmatrix},\mathbf{k}=\begin{bmatrix}\mathbf{k}_{1}\\ \vdots\\ \mathbf{k}_{K}\end{bmatrix},\mathbf{u}=\begin{bmatrix}\mathbf{u}_{1}\\ \vdots\\ \mathbf{u}_{K}\end{bmatrix}$.
Note that while the closed form solution requires the inversion of a $3K\times 3K$ matrix, the number of interaction points, $K$ is often small, leading to a very fast solution for Kelvinlet. The total Kelvinlet deformation model, including the analytic solution to the optimization, can be performed in less than 1 ms for two simultaneously engaged graspers (i.e., $K\approx 20$) using native GPU acceleration in PyTorch on an NVIDIA A6000 GPU and on a mesh with $\sim 10\text{k}$ nodes.
Having established a fast physics-based solution, we now focus on integrating Kelvinlets into our models.
3.2 Method 1: Kelvinlet-Based Residual Learning
For an initial displacement $\mathbf{u}_{s}$ located at $\mathbf{x}_{s}$, let $\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ denote the ground-truth displacement field. Instead of directly regressing $\mathbf{u}_{\text{true}}$, we introduce a residual learning framework where a neural network estimates the deviation from the Kelvinlet prior, which serves as an initial approximation. Specifically, we define the residual displacement as $\mathbf{r}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s})=\mathbf{u}_{\text{true}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s})-\mathbf{u}_{\varepsilon}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s}),$ where $\mathbf{u}_{\varepsilon}$ represents the Kelvinlet solution. The neural network, parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{d}$, learns to approximate this residual via $\hat{\mathbf{r}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta}).$ The corrected displacement field is then given by $\tilde{\mathbf{u}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})=\mathbf{u}_{\varepsilon}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s})+\hat{\mathbf{r}}(\mathbf{x};\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta}).$ Let $\|\mathbf{u}\|_{L_{2}(\Omega)}^{2}=\int_{\Omega}\|\mathbf{u}(\mathbf{x})\|^{2}d\mathbf{x}$, then to train the network, we minimize the loss:
$\mathcal{L}_{\text{residual}}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\Big[\|\mathbf{u}_{\text{true}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})\\ -\tilde{\mathbf{u}}(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})\|^{2}_{L_{2}(\Omega)}\Big],$ (11)
where $p$ represents the distribution of all possible pointwise interactions with the tissue, and $\Omega$ denotes the tissue domain. By focusing on learning only the correction beyond the Kelvinlet solution, this approach leverages the physics priors directly into its formulation.
3.3 Method 2: Kelvinlets as Regularization
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]\\ +\lambda_{\text{reg}}\,\mathbb{E}_{(\mathbf{x}^{\prime}_{s},\mathbf{u}^{\prime}_{s})\sim q}\left[\left\|\mathbf{u}_{\varepsilon}-\hat{\mathbf{u}}\right\|^{2}_{L_{2}(\Omega)}\right]$ (12)
where $\lambda_{\text{reg}}>0$ is the regularization coefficient. For clarity and brevity, we omit the dependencies $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s})$ and $(\cdot;\mathbf{x}_{s},\mathbf{u}_{s},\boldsymbol{\theta})$ in the notation. Note that we explicitly separate $(\mathbf{x}_{s},\mathbf{u}_{s})$ from $(\mathbf{x}_{s}^{\prime},\mathbf{u}^{\prime}_{s})$ because, in practice, the two distributions $p$ and $q$ could differ, and also while Monte Carlo integration is used for both terms, the first term requires FEM simulations to obtain $\mathbf{u}_{\text{true}}$ (which is computationally expensive), whereas the second term only involves Kelvinlet evaluations (much faster). We will describe the distributions $p$ and $q$ used in our experiments in the following section.
Figure 2 illustrates the proposed residual- and regularization-based learning methods, along with the displacement norm visualizations for all methods on a representative sample.
@@ 4 Experiments
Figure 3: Qualitative comparison of deformation predictions for the best (PointNet) and worst (GraphGPS) models in linear regimes.
@@
We first evaluate several deformation-specific baselines designed for soft tissue modeling. PhysGNN [33] is a mesh-based GNN trained on FEM simulations; we enhance it with a global attention layer after each graph convolution to mitigate the well-known oversmoothing associated with GNNs [35]. PEGNN [32] models contact dynamics via cross-graph attention; we retain only its soft-tissue encoder to isolate local geometric features. For SeeSaw [9], we reimplement the method and adapt it to a point-cloud-based simulation setup. From the diffeomorphic mapping approach [12], we extract only the deformation network to enable direct supervision. We also include a Transformer-based KAN [16], limited to single-step prediction for fair comparison. To assess architecture-agnostic effects of Kelvinlet priors, we additionally evaluate U-Net [22], SAB [19], GraphGPS [29], and PointNet [28]. U-Net is converted into a DeepSet by replacing convolutions with shared MLPs. SAB uses two self-attention layers and an MLP head. GraphGPS and PointNet are simplified by reducing their depth to prevent overfitting on high-resolution meshes.
@@ 4.1 Results and Analysis
Table 1: Deformation Capture Mean (DCM) (%) on individual graspers in linear simulations for modified baselines and our method, with and without Kelvinlet priors.
Ours
Model Base +Res. +Reg.
PhysGNN[33] 90.12 (0.13) 91.77 (0.11) 92.65 (0.10)
SeeSaw[9] 89.24 (0.15) 91.02 (0.10) 91.83 (0.08)
Physics-Encoded GNN[32] 88.55 (0.14) 90.44 (0.09) 91.58 (0.12)
Transformer-KAN[16] 90.89 (0.12) 92.01 (0.08) 93.22 (0.07)
DiffoMorphic Deformation[12] 91.14 (0.10) 92.88 (0.11) 94.33 (0.09)
PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09)
U-Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11)
GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23)
SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04)
Kelvinlet 65.24 (0) - -
Average Gain +1.23 +2.39
Minimum/Maximum Gain +0.43/1.89 +1.01/3.19
Table 2: Inference time (ms) on an NVIDIA RTX A6000 GPU for a single grasper with individual models in linear simulations, with and without Kelvinlet priors.
Inference time (ms) Ours
Model Base +Res. +Reg.
PhysGNN[33] 1.71 (0.11) 1.84 (0.13) 1.71 (0.11)
SeeSaw[9] 1.56 (0.09) 1.69 (0.08) 1.56 (0.09)
Physics-Encoded GNN[32] 1.65 (0.08) 1.79 (0.08) 1.65 (0.08)
Transformer-KAN[16] 2.22 (0.17) 2.32 (0.16) 2.22 (0.17)
DiffoMorphic Deformation[12] 2.16 (0.12) 2.35 (0.11) 2.16 (0.12)
PointNet[28] 1.16 (0.10) 1.23 (0.13) 1.16 (0.10)
U-Net[22] 1.55 (0.08) 1.64 (0.07) 1.55 (0.08)
GraphGPS[29] 2.60 (0.15) 2.78 (0.15) 2.60 (0.15)
SAB[19] 2.30 (0.02) 2.44 (0.21) 2.30 (0.20)
Kelvinlet 0.15 (0.052)
Table 3: Performance of various models on linear simulations with individual graspers, combined graspers, and multitask learning. DCM (%) is reported for single-task and multi-task settings with standard deviations.
Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
Base +Res. +Reg. Base +Res. +Reg. Base +Res. +Reg.
U -Net[22] 92.11 (0.12) 93.56 (0.10) 94.04 (0.11) 88.23 (0.08) 89.67 (0.09) 90.01 (0.08) 91.20 (0.09) 92.33 (0.10) 92.88 (0.08)
GraphGPS[29] 88.60 (0.21) 89.45 (0.15) 89.61 (0.23) 87.33 (0.06) 87.92 (0.02) 88.01 (0.04) 88.35 (0.13) 88.91 (0.10) 89.45 (0.09)
PointNet[28] 94.23 (0.11) 94.66 (0.09) 96.54 (0.09) 89.79 (0.07) 90.11 (0.07) 91.22 (0.07) 94.00 (0.08) 94.12 (0.07) 95.01 (0.06)
SAB[19] 93.42 (0.11) 94.60 (0.04) 96.03 (0.04) 88.05 (0.09) 89.91 (0.11) 91.02 (0.04) 92.54 (0.06) 93.66 (0.07) 94.42 (0.06)
Kelvinlet 65.24 (0) 64.33 (0) 64.82 (0)
Table 4: Performance of various models on nonlinear simulations with individual graspers, combined graspers, and multitask learning. DCM (%) is reported for single-task and multi-task settings with standard deviations.
Model Indiv. Graspers (DCM) Comb. Graspers (DCM) Multitask (DCM)
Base +Res. +Reg. Base +Res. +Reg. Base +Res. +Reg.
U -Net[22] 80.56 (0.10) 84.22 (0.09) 84.88 (0.11) 76.91 (0.05) 79.83 (0.08) 82.10 (0.21) 78.10 (0.12) 80.45 (0.09) 81.33 (0.14)
GraphGPS[29] 77.52 (0.17) 80.63 (0.13) 82.11 (0.09) 73.72 (0.15) 77.92 (0.10) 80.66 (0.16) 76.24 (0.18) 79.31 (0.15) 80.42 (0.10)
PointNet[28] 81.25 (0.22) 84.65 (0.19) 85.05 (0.15) 79.56 (0.18) 82.66 (0.16) 83.33 (0.11) 83.25 (0.17) 84.75 (0.13) 85.21 (0.09)
SAB[19] 80.75 (0.19) 84.44 (0.07) 84.75 (0.10) 78.88 (0.11) 81.50 (0.15) 82.83 (0.16) 81.22 (0.10) 83.01 (0.08) 84.00 (0.07)
Kelvinlet 58.88 (0) 51.67 (0) 55.56 (0)
To quantitatively assess model performance, we utilize the Deformation Capture Mean (DCM) metric, which measures how well a predicted deformation field approximates the ground-truth displacement field. The DCM is computed as:
@@
Table 1 shows that integrating Kelvinlet-based priors improves deformation accuracy across all models. Residual learning yields an average gain of +1.23%, while Kelvinlet regularization achieves a stronger improvement of +2.39%, with the highest gain of +3.19% observed in DiffoMorphic Deformation. PointNet benefits the most in absolute performance, reaching 96.54% DCM. The analytical Kelvinlet alone performs poorly (65.24%), underscoring the value of learning residuals.
As seen in Table 2, our methods add minimal runtime overhead. Regularized variants match the base inference time, while residual versions increase latency by only $\sim$0.1ms. Kelvinlet inference alone is extremely fast (0.15ms) but not accurate enough, justifying its role as a prior. Overall, our approach enhances accuracy without compromising efficiency.
Tables 3 and 4 report DCMs across models and settings. Each experiment runs with three random seeds, reporting means and standard deviations. Models perform better on linear cases, as nonlinear deformations introduce gravity effects, non-proportional stress-strain relationships, and greater sensitivity to perturbations. Notably, Kelvinlet priors consistently improve DCMs, despite being derived for linear elasticity, showing their effectiveness even in nonlinear regimes.
@@ 5 Conclusion
Acknowledgments
This project is supported by the Wellcome Leap Surgery: Assess/Validate/Expand (SAVE) program.